What volume of .06892 M NaOH is required to react completely with 41.25-mLof .06053 M H2SO4?
1
Expert's answer
2012-10-17T11:16:05-0400
n of H2SO4 in 41.25 ml is equal to
0.06053 mol in 1000 ml

n of acid
=0.002497 mol

H2SO4 + 2NaOH=Na2SO4 + H2O

n of NaOH =2*n of H2SO4 =
0.004994 mol

0.004994 mol in x ml
0.06892 mol in 1000
ml

x=72.46 ml